OpenStudy (anonymous):

Can someone help me with Geometry please ? I don't need answers , I just need help understanding .

OpenStudy (oleg3321):

no. you cuss too much

OpenStudy (anonymous):

okay ? thats cool . like I said , if your not going to help me then just don't comment please !

OpenStudy (anonymous):

Fill in each missing step. Given: mÐPQR = x– 5, mÐSQR = x – 11, and mÐPQS = 100. Find x.

OpenStudy (anonymous):

Drawing not to scale. mÐPQR +mÐSQR = mÐPQS a. __________ x– 5 + x– 11 = 100 b. Substitution Property 2x– 16 = 100 c. Simplify 2x = 116 d. __________ x = 58 e. Division Property of Equality Select one: a. Angle Addition Postulate; Subtraction Property of Equality b. Protractor Postulate; Addition Property of Equality c. Angle Addition Postulate; Addition Property of Equality d. Protractor Postulate; Subtraction Property of Equality